# DFT-s-OFDM-based On-Off Keying for Low-Power Wake-Up Signal

### v1

@@ III-A4 FD Post-Processing
where $(N_{\rm sc}-N_{\rm pulse})$ is the spectrum extension (SE), $L$ is an integer shift, $W[k]$ is an FDSS window, and $\eta$ is a normalization factor.
Conventional DFT-s-OFDM is with $N_{\rm pulse}=N_{\rm sc}$, $L=0$, and then $X^{W}[k]=\eta D[k]$. SE and FDSS have been considered for low-PAPR (Peak-to-Average Power Ratio) benefit [16], and thus for controlling the envelope of the signal, at the cost of breaking multi-carrier waveform orthogonality and creating ICI. These methods provide similar degrees of freedom in shaping the OOK waveform while for LP-WUS there is no obvious cost in breaking orthogonality.
In addition, SE here enables control of $N_{\rm pulse}\leq N_{\rm sc}$ such that the bit spreading factor $N_{\rm seg}=N_{\rm pulse}/N_{\rm bit}$ can be set to be an integer. For example for transmitting $N_{\rm bit}=8$ bits in a bandwidth of $N_{\rm sc}=12\times 11=132$ then $N_{\rm bit}/N_{\rm sc}$ is not an integer; but by selecting $N_{\rm pulse}=128$ one is able to satisfy $N_{\rm seg}=N_{\rm bit}/N_{\rm pulse}$ integer.
As for FDSS, we consider possibly complex window but limited to the form $W[k]=e^{-\frac{{\rm j}2\pi}{N_{\rm fft}}T_{\rm shift}k}W_{R}[k]$ where $W_{R}[k]$ is a real and symmetric FDSS window and $T_{\rm shift}$ corresponds to a cyclic time shift of the waveform. In this paper we use a Kaiser window, known to concentrate well the energy of DFT-s-OFDM pulses [17], and defined as $W_{R}[k]=I_{0}\left(\beta\sqrt{1-\frac{(k-\gamma)^{2}}{\gamma^{2}}}\right)/I_{0}\left(\beta\right)$ with shaping parameter $\beta$ and where $\gamma=(N_{\rm sc}-1)/2$ and $I_{0}(\cdot)$ represents the zeroth-order modified Bessel function of the first kind. With $\beta=0$ $W_{R}[k]=1$ for all $k$ i.e. there is no FDSS; and increasing the shaping parameter $\beta$ increases the FD energy concentration.
The normalization factor $\eta$ adjust the LP-WUS power. In [2], LP-WUS power spectral density should take into account the guard band usage to be aligned with the concurrent average symbol power $P_{s}$, i.e., $\eta$ in (7) is selected satisfying $\sum_{k=0}^{N_{\rm sc}-1}\operatorname{E}\left[|X^{W}[k]|^{2}\right]=P_{s}(N_{\rm sc}+N_{\rm GB})$.
Figure 2: Illustration of DFT-s-OFDM based OOK modulation
(a) $N_{\rm pulse}=12$, plain DFT-s-OFDM
@@ III-B TD Pulse Multiplexing Aspect
$s^{W}[n]=\eta\sum_{p=0}^{N_{\rm bit}-1}b[p]O_{p}[n]$ (12)
where $O_{p}[n]=\sum_{m=pN_{\rm seg}}^{(p+1)N_{\rm seg}-1}r[m]g_{m}[n]$ is the OOK symbol for the $p$th bit. Each pulse $g_{m}[n]$ has most of its energy confined in the sample interval $\left[m\frac{N_{\rm fft}}{N_{\rm pulse}},(m+1)\frac{N_{\rm fft}}{N_{\rm pulse}}\right]$ with an energy peak in the middle. Most of the potential combined energy of $O_{p}[n]$ is thus in the sample interval $\left[\frac{N_{\rm fft}}{N_{\rm bit}}p,\ldots,\frac{N_{\rm fft}}{N_{\rm bit}}(p+1)\right]$. Fig. 3 illustrates this pulse modulation aspect within one OFDM symbol. Remark in Fig. 3(a) that the first pulse is always peaking at time zero with its energy spread equally at the beginning and end of the OFDM symbol. This will result in a circular leakage from the first OOK symbol to the last OOK symbol, but can be attenuated by shifting all pulses to the right using a linear phase in $W[k]$ of $T_{\rm shift}=\frac{N_{\rm fft}}{2N_{\rm pulse}}$ as in Fig. 3(b). Also in Fig. 3(b), SE is applied, which decreases the number of pulses and spread them apart; and the pulse kernel (10) is shaped with $\beta=2$ which has the typically effect to enlarge its main lobe while attenuating it sidelobes. As a result of both SE and FDSS the pulses are non-orthogonal.
@@ III-C Simplified Subcarrier Coefficients
Consider as bit-spreading a common overlaid sequence $r_{0}[m]$ possibly covered by linear phase ramp as
$r[l]=e^{{\rm j}\Phi l}r_{0}[m-N_{\rm seg}l_{m}].$ (13)
Figure 4: Equivalent FD implementation of DFT-s-OFDM based OOK in Fig. 2 when bit-spreading is made of common overlaid sequence $\tilde{r}_{l}[n]=r_{0}[n]$.
@@ III-C1 Two-Layer FD Sequences
where $B(f)=\sum_{m=0}^{N_{\rm bit}-1}b[m]e^{-{\rm j}\frac{2\pi}{N_{\rm bit}}fm}$ is the interpolated DFT [^2] Analogous to the discrete-time Fourier transform (DTFT) of the bits; and $R_{0}(f)=\sum_{l=0}^{N_{\rm seg}-1}r_{0}[l]e^{-{\rm j}\frac{2\pi}{N_{\rm pulse}}fl}$ interpolated DFT of the overlaid sequence $r_{0}[n]$
@@
$\mathcal{B}[k]=\sum_{m=0}^{N_{\rm bit}-1}b[m]e^{-{\rm j}\frac{2\pi}{N_{\rm bit}}km}$ (16)
and where the FD overlaid sequence $R_{0}[k]=\sum_{l=0}^{N_{\rm seg}-1}r_{0}[l]e^{-{\rm j}\frac{2\pi}{N_{\rm pulse}}kl}$ is the $N_{\rm pulse}$-point DFT of the overlaid sequence $r_{0}[n]$ with zero padding.
Lem. 1 and Corr. 1 provides a different interpretation of DFT-s-OFDM-based OOK. The sequence $R_{0}$ provides a common shaping window, constructing a ON symbol kernel, for all possible data-dependent sequences $B$. This ON kernel does not depends of the bits values but still depends on the bit number via $N_{\rm seg}$, setting a proper ON symbol duration. The sequence $B$ carries the bit information by superposing multiple time-shifted version of the same ON symbol kernel. This is illustrated in Fig. 5 for the bit string $(0,1,0,1)$, with $N_{\rm sc}=72$, a Zadoff-Chu (ZC) sequence as overlaid sequence, and no post-processing. As shown in the figure, the IFFT of $R_{0}[k]$ creates a single ON symbol while the IFFT of $B[k]$ generates sinc-pulses at different time instants. and the resulting OFDM signal is the convolution of these two signals, up to a scaling.
Figure 5: Illustration of the layered FD sequences $B$ and $R_{0}$.
@@ IV-A Flattening ON Symbols in DFT-s-OFDM-based OOK
With a phase ramp as bit-spreading, the WUS subcarrier coefficients (7) can be further simplified using Lem. 1 and simplifying $R_{0}(f)$ with the sum exponential formula.
@@
$\!\!X^{W}[k]=\eta W[k]R_{0}\left(k-\textstyle\frac{N_{\rm pulse}\Phi}{2\pi}+L\right)B\left(k-\textstyle\frac{N_{\rm pulse}\Phi}{2\pi}+L\right)$ (19)
@@
in which $\alpha=e^{{\rm j}\pi(\frac{1}{N_{\rm bit}}-\frac{1}{N_{\rm pulse}})\left(\frac{\Phi N_{\rm pulse}}{2\pi}-L\right)}$.
The global phase $\alpha$ has no impact on the signal envelope and can be disregarded.
@@ IV-B Rectangular Zero-DC OOK
Due to the sinc-like function $R_{0}(f)$, the power distribution of the subcarriers coefficients will be concentrated in a dominant lobe whose position is also shifted by the phase ramp. To ensure that the LP-WUS average spectrum is symmetrically distributed around the DC, different values of $\lambda$ can be selected, or the sign of $\Phi$ can be alternated in each OFDM symbol. Fig. 7 illustrates the resulting DC nulling effect with $N_{\rm pulse}=N_{\rm sc}=48$, $N_{\rm bit}=8$, $k_{\rm null}=24$, and using $\Phi=\pm\frac{2}{3}\pi$ with alternating sign. As shown, the middle subcarrier has been canceled, while the OOK waveform retains a rectangular shape since $\Phi=\frac{2}{3}$ is close enough to $\pi$.
@@ IV-C Least-Squares (LS) Waveform
In 3GPP LP-WUS study [2] as well as in the on-going 3GPP Ambient-IoT study [21], one of the main approaches is from [14], which provides the LS approximation to an ideal rectangular OOK signal. The subcarrier coefficients are obtained using a so-called “LS precoding” [2], which differs from DFT-precoding in (6). In fact, this method essentially involves a larger DFT of size matching the OFDM IFFT, followed by a truncation to match the bandwidth. See Appendix -D for further details.
The LS subcarrier coefficients following [14] can be written for $k=0,\ldots,N_{\rm sc}-1$ as
@@
where for $l=0,\ldots,N_{\rm fft}-1$
$D^{\rm LS}[l]=\sum_{n=0}^{N_{\rm fft}-1}b_{s}^{\rm LS}[n]e^{-j\frac{2\pi}{N_{\rm fft}}nl}$ (23)
is a $N_{\rm fft}$-point DFT of the spreaded-bit sequence $b_{s}^{\rm LS}$, obtained by repeating the bit string $b[m]$ of length $N_{\rm bit}$ with spreading factor $N_{\rm seg}^{\rm LS}=N_{\rm fft}/N_{\rm bit}$ (assumed being an integer). Since the perfectly-rectangular OOK signal $b_{s}^{\rm LS}[k]$ is real, its spectrum is symmetrically concentrated around the zero frequency. Therefore the LS approximation to $b_{s}^{\rm LS}[k]$ with constrained bandwidth is given (22) which preserves the dominant subcarrier coefficients around the zero frequency.
The subcarrier coefficients (22) can be further simplified.
@@
The proof is very similar to Lem. 1 and Corr. 2, and therefore is omitted. By identifying with Corr. 2, one derives
@@
Finally, we remark that in narrowband scenarios such as those in Ambient-IoT applications where larger bit rate and smaller bandwidth than LP-WUS are considered, approximating an ideal –inherently wideband– rectangular signal becomes difficult. In such cases, investigating alternative DFT-s-OFDM-based designs may offer interesting solutions. We provide an example in Fig. 9 for $N_{\rm bit}=6$ bits per OFDM symbol using $N_{\rm sc}=12$ subcarriers, as considered in [21]. This figure illustrates the DFT-s-OFDM-based implementation therein, corresponding here to $N_{\rm pulse}=N_{\rm sc}$, $L=N_{\rm sc}/2$ (referred as ‘fftshift’), and $\Phi=0$. This implementation satisfies the relationship between $L$ and $\Phi$ as in Cor. 3 but differs from the LS waveform by not including the FDSS (26), which results primarily in a time shifting difference. Fig. 9 also shows an alternative signal design using the minimum number of DFT-s-OFDM pulses $N_{\rm pulse}=N_{\rm bit}$ combined with more spectrum shaping using a Kaiser window with $\beta=4$; and both the $(\Phi,L)$ relationship from Cor. 3 and from (18). Since $N_{\rm sc}$ is even and small, we have $\Phi=\pi\left(L/3+2\right)$ in Cor. 3, which differs from $\Phi=\pi\left(L/3+11/6\right)$ in (18). As shown, the DFT-s-OFDM waveform with $N_{\rm pulse}=N_{\rm bit}$ and $\Phi$ from (18) provides better pulse transition and smaller sidelobe in OFF symbols, resulting in power-boosted ON symbols which could facilitate detection.
@@ V Shaped OOK for 5G LP-WUS
In 3GPP LP-WUS study [2], only $N_{\rm bit}=1$, $2$, or $4$ are considered, while the ADC sampling rate at the LR is assumed to be much higher than such low OOK data rates. Indeed, it was shown therein that an ADC sampling rate of, for example, 3.84 MHz was usable to improve performance while having little impact on the LR’s power consumption. With sufficient ADC precision at LR, the OOK detection performance principally depends on the energy split between the ON and OFF symbols, while minimizing the envelope fluctuation inside ON symbols may actually matter little. Therefore, a rectangular-like OOK waveform as targeted in the previous section does not provide the best performance for 5G LP-WUS, and in this section, we discuss two signal features that were recognized in [2] as useful for performance improvement.
V-A Spreading OOK Power Spectrum
…for improving performance against frequency-selective fading: An LP-WUS bandwidth no larger than $5$ MHz was recommended in [2] to facilitate deployment, amounting to less than 144 subcarriers with $30$ kHz spacing in 3GPP numerology. At the same time, larger signal bandwidth can increase frequency diversity and thus robustness against frequency selective fading, such that most sources in [2] considered exploiting as much as possible of this $5$ MHz. In this context of a relatively large bandwidth with a low data rate, rectangular-like OOK signals discussed in the previous section have concentrated power spectrum that makes them vulnerable to frequency-selectivity of wireless fading. Alternative signal designs with more spread-out spectrum, but without compromising much on the ON/OFF energy split, can thus improve performance [^4] However, for the purpose of designing OOK synchronization signal, an OOK signal with narrow spectrum can provide benefits for performing frequency-synchronization in frequency-selective channel [22].. Below are methods to achieve this.

### v2

@@ III-A4 FD Post-Processing
where $(N_{\rm sc}-N_{\rm pulse})$ is the spectrum extension (SE), $L$ is an integer shift, $W[k]$ is an FDSS window, and $\eta$ is a normalization factor. Conventional DFT-s-OFDM is with $N_{\rm pulse}=N_{\rm sc}$, $L=0$, and then $X^{W}[k]=\eta D[k]$.
SE enables to control the number of pulses $N_{\rm pulse}\leq N_{\rm sc}$ such that the spreading factor $N_{\rm seg}=N_{\rm pulse}/N_{\rm bit}$ is an integer. For example, with $N_{\rm bit}=8$ and bandwidth $N_{\rm sc}=12\times 11=132$, $N_{\rm bit}/N_{\rm sc}$ is not an integer; so instead by selecting $N_{\rm pulse}=128$ one gets an integer spreading factor $N_{\rm seg}=N_{\rm pulse}/N_{\rm bit}$.
FDSS enables shaping of the ON symbols. We consider an FDSS window limited to the form $W[k]=e^{-\frac{{\rm j}2\pi}{N_{\rm fft}}T_{\rm shift}k}W_{R}[k]$ where $W_{R}[k]$ is a real and symmetric FDSS window and $T_{\rm shift}$ corresponds to a cyclic time shift of the waveform. For numerical evaluations, because it concentrates well the energy of DFT-s-OFDM pulses [25], we will use a Kaiser window defined as $W_{R}[k]=I_{0}\left(\beta\sqrt{1-\frac{(k-\gamma)^{2}}{\gamma^{2}}}\right)/I_{0}\left(\beta\right)$ with shaping parameter $\beta$ and where $\gamma=(N_{\rm sc}-1)/2$ and $I_{0}(\cdot)$ represents the zeroth-order modified Bessel function of the first kind. With $\beta=0$, $W_{R}[k]=1$ for all $k$ i.e. there is no shaping; and as the parameter $\beta$ increases, the FD energy concentration increases.
The normalization factor $\eta$ adjust the LP-WUS power. In [2], LP-WUS power spectral density should take into account the guard band usage to be aligned with the concurrent average symbol power $P_{s}$, i.e., $\eta$ in (7) is selected satisfying $\sum_{k=0}^{N_{\rm sc}-1}\operatorname{E}\left[|X^{W}[k]|^{2}\right]=P_{s}(N_{\rm sc}+2N_{\rm GB})$.
@@ III-B TD Pulse Multiplexing Aspect
$s^{W}[n]=\eta\sum_{l=0}^{N_{\rm bit}-1}b[l]O_{l}[n]$ (12)
where
$O_{l}[n]=\!\!\!\sum_{m=lN_{\rm seg}}^{(l+1)N_{\rm seg}-1}\!\!\!\!\!\!r[m]g_{m}[n]=\sum_{m=0}^{N_{\rm seg}-1}\!\!\!\tilde{r}_{l}[m]g_{m+lN_{\rm seg}}[n]$ (13)
is the OOK symbol for the $l$th bit. Each pulse $g_{m}[n]$ has most of its energy confined in the sample interval $n\in\left[m\frac{N_{\rm fft}}{N_{\rm pulse}},(m+1)\frac{N_{\rm fft}}{N_{\rm pulse}}\right]$ with an energy peak in the middle. Hence, most of the potential combined energy of $O_{l}[n]$ is in the sample interval $n\in\left[\frac{N_{\rm fft}}{N_{\rm bit}}l,\ldots,\frac{N_{\rm fft}}{N_{\rm bit}}(l+1)\right]$ suitable for creating an ON symbol in it, while combined energy leakage from these pulses in other sample intervals can contribute to undesired signal level in OFF symbols. Fig. 3 illustrates this pulse modulation aspect within one OFDM symbol. Remark in Fig. 3(a) that the first pulse is always peaking at time zero with its energy spread equally at the beginning and end of the OFDM symbol. This will result in a circular leakage from the first OOK symbol to the last OOK symbol, but can be attenuated by shifting all pulses to the right using a linear phase in $W[k]$ of $T_{\rm shift}=\frac{N_{\rm fft}}{2N_{\rm pulse}}$ as in Fig. 3(b). Also in Fig. 3(b), SE is applied, which decreases the number of pulses and spread them apart; and the pulse kernel (10) is shaped with $\beta=2$ which has the typically effect to enlarge its main lobe while attenuating it sidelobes. As a result of both SE and FDSS the pulses are non-orthogonal.
(a) $N_{\rm pulse}=12$, plain DFT-s-OFDM
@@ III-C Simplified Subcarrier Coefficients
Consider as bit-spreading a common overlaid sequence $r_{0}[n]$ possibly covered by linear phase ramp as
$r[m]=e^{{\rm j}\Phi m}r_{0}[m\;{\rm mod}\;N_{\rm seg}].$ (14)
Figure 4: Equivalent FD implementation of DFT-s-OFDM-based OOK in Fig. 1 when bit-spreading is made of common overlaid sequence $\tilde{r}_{l}[n]=r_{0}[n]$.
@@ III-C1 Two-Layer FD Sequences
where $B(f)=\sum_{l=0}^{N_{\rm bit}-1}b[l]e^{-{\rm j}\frac{2\pi}{N_{\rm bit}}fl}$ is the interpolated DFT [^2] Analogous to the discrete-time Fourier transform (DTFT) of the bits; and $R_{0}(f)=\sum_{m=0}^{N_{\rm seg}-1}r_{0}[m]e^{-{\rm j}\frac{2\pi}{N_{\rm pulse}}fm}$ is the interpolated DFT of the overlaid sequence $r_{0}[n]$
@@
$\mathcal{B}[k]=\sum_{l=0}^{N_{\rm bit}-1}b[l]e^{-{\rm j}\frac{2\pi}{N_{\rm bit}}kl}$ (17)
and where the FD overlaid sequence $R_{0}[k]=\sum_{m=0}^{N_{\rm seg}-1}r_{0}[m]e^{-{\rm j}\frac{2\pi}{N_{\rm pulse}}km}$ is the $N_{\rm pulse}$-point DFT of the overlaid sequence $r_{0}[n]$ with zero padding.
Lem. 1 and Corr. 1 provides a different interpretation of DFT-s-OFDM-based OOK. The sequence $R_{0}[k]$, along with FDSS-SE, provides a common shaping window, constructing an OOK symbol kernel, for all possible data-dependent sequences $B[k]$. This OOK kernel does not depends of the bits values but still depends on the bit number via $N_{\rm seg}$, setting a proper ON symbol duration. The sequence $B[k]$ carries the bit information by superposing multiple time-shifted version of the same OOK symbol kernel. Concretely, with (14) and if $N_{\rm fft}/N_{\rm bit}$ is an integer, the OOK symbols as given in (13) become a circular shift of the OOK symbol kernel $O_{0}[n]$, up to a phase coefficient, as
$O_{l}[n]=e^{-{\rm j}\frac{2\pi l}{N_{\rm bit}}\left(L-\frac{\Phi N_{\rm pulse}}{2\pi}\right)}O_{0}\left[n-l\frac{N_{\rm fft}}{N_{\rm bit}}\right].$ (18)
This is illustrated in Fig. 5 for the bit string $(0,1,0,1)$, with $N_{\rm sc}=72$, a ZC overlaid sequence, and no post-processing. As shown in the figure, the IFFT of $R_{0}[k]$ creates a single ON symbol while the IFFT of $B[k]$ generates sinc-pulses at different time instants, and the resulting OFDM signal is the convolution of these two signals, up to a scaling.
Figure 5: Illustration of the layered FD sequences $B[k]$ and $R_{0}[k]$.
@@ IV-A Flattening ON Symbols in DFT-s-OFDM-based OOK
With a phase ramp as bit-spreading, the WUS subcarrier coefficients (7) can be further simplified as follows (see Appendix -A).
@@
$\!\!X^{W}[k]=\eta W[k]R_{0}\left(k-\textstyle\frac{\Phi N_{\rm pulse}}{2\pi}+L\right)B\left(k-\textstyle\frac{\Phi N_{\rm pulse}}{2\pi}+L\right)$ (21)
@@
in which [^4] The global phase $\alpha$ has no impact on the signal envelope and can be disregarded. $\alpha=e^{{\rm j}\pi(\frac{1}{N_{\rm bit}}-\frac{1}{N_{\rm pulse}})\left(\frac{\Phi N_{\rm pulse}}{2\pi}-L\right)}$.
@@ IV-B Rectangular Zero-DC OOK
Due to the sinc-like function $R_{0}(f)$, the power distribution of the subcarriers coefficients will be concentrated in a dominant lobe whose position is also shifted by the phase ramp. To ensure that the LP-WUS average spectrum is symmetrically distributed around the DC, different values of $\lambda$ can be selected, or the sign of $\Phi$ can be alternated in each OFDM symbol. Fig. 7 illustrates the resulting DC nulling effect with $N_{\rm pulse}=N_{\rm sc}=48$, $N_{\rm bit}=8$, $k_{\rm null}=24$, and using $\Phi=\pm\frac{2}{3}\pi$ with alternating sign. As shown, the middle subcarrier has been canceled, while the OOK waveform retains a rectangular shape since $\Phi=\frac{2}{3}\pi$ is close enough to $\pi$.
@@ IV-C Least-Squares (LS) Waveform
In 3GPP LP-WUS and A-IoT studies, one of the main approaches was from [14] which provides the LS approximation to an ideal rectangular OOK signal. The corresponding subcarrier coefficients are obtained using a so-called “LS precoding” in [2] which is essentially a larger DFT of size matching the OFDM IFFT, followed by a truncation to match the subcarrier allocation, see Appendix -D for further details.
The LS subcarrier coefficients obtained in [14] can be written for $k=0,\ldots,N_{\rm sc}-1$ as
@@
where for $k^{\prime}=0,\ldots,N_{\rm fft}-1$
$D^{\rm LS}[k^{\prime}]=\sum_{n=0}^{N_{\rm fft}-1}b_{\rm r}^{\rm LS}[n]e^{-j\frac{2\pi}{N_{\rm fft}}nk^{\prime}}$ (25)
is a $N_{\rm fft}$-point DFT of the spread-bit sequence $b_{\rm r}^{\rm LS}$ obtained by repeating the bit string $b[m]$ of length $N_{\rm bit}$ with spreading factor $N_{\rm seg}^{\rm LS}=N_{\rm fft}/N_{\rm bit}$ (assumed being an integer). Since the perfectly-rectangular OOK signal $b_{\rm r}^{\rm LS}[k]$ is real, its spectrum is symmetrically concentrated around the zero frequency. Therefore the LS approximation to $b_{\rm r}^{\rm LS}[k]$ with constrained bandwidth is given by (24) which preserves the dominant subcarrier coefficients around the zero frequency.
The LS subcarrier coefficients (24) can be further simplified (see Appendix. -A).
@@
By identification with Corr. 2, it follows
@@
Finally, we remark that in narrowband scenarios such as those in A-IoT applications where larger bit rate and smaller bandwidth than LP-WUS are considered, approximating an ideal –inherently wideband– rectangular signal becomes difficult. In such cases, investigating alternative DFT-s-OFDM-based designs may offer interesting solutions. We provide an example in Fig. 9 for $N_{\rm bit}=6$ bits per OFDM symbol using $N_{\rm sc}=12$ subcarriers, as considered in [29]. This figure illustrates the DFT-s-OFDM-based implementation therein, corresponding here to $N_{\rm pulse}=N_{\rm sc}$, $L=N_{\rm sc}/2$ (referred as ‘fftshift’), and $\Phi=0$. This implementation satisfies the relationship between $L$ and $\Phi$ as in Cor. 3 but differs from the LS waveform by not including the FDSS (28), which results primarily in a time shifting difference. Fig. 9 also shows an alternative signal design using the minimum number of DFT-s-OFDM pulses $N_{\rm pulse}=N_{\rm bit}$ combined with more spectrum shaping using a Kaiser window with $\beta=4$; and both the $(\Phi,L)$ relationship from Cor. 3 and from (20). Since $N_{\rm sc}$ is even and small, we have $\Phi=\pi\left(L/3+2\right)$ in Cor. 3, which differs from $\Phi=\pi\left(L/3+11/6\right)$ in (20). As shown, the DFT-s-OFDM waveform with $N_{\rm pulse}=N_{\rm bit}$ and $\Phi$ from (20) provides better pulse transition and smaller sidelobe in OFF symbols, resulting in power-boosted ON symbols which could facilitate detection.
@@ V Shaped OOK for 5G LP-WUS
In 5G LP-WUS, only $N_{\rm bit}=1$, $2$, or $4$ can be transmitted, while the ADC sampling rate at the LR is assumed to be much higher than such low OOK data rates. Indeed, it was shown in [2] that an ADC sampling rate of, for example, 3.84 MHz was usable to improve performance while having little impact on the LR’s power consumption. With sufficient ADC precision at LR, the OOK detection performance principally depends on the energy split between the ON and OFF symbols, while minimizing the envelope fluctuation inside ON symbols may actually matter little. Therefore, the standard rectangular-like OOK waveform as targeted in the previous section does not provide the best performance for LP-WUS, and in this section, we propose and analyze FD and TD shaping features for performance improvement in such application environments.
V-A FD Shaping for Improved Performance Against Frequency-Selective Fading
An LP-WUS bandwidth no larger than $5$ MHz was recommended in [2] to facilitate deployment, amounting to less than 144 subcarriers with $30$ kHz spacing in 3GPP numerology. At the same time, larger signal bandwidth can increase frequency diversity and thus robustness against frequency selective fading, such that most sources in [2] considered exploiting as much as possible of this $5$ MHz. In this context of a relatively large bandwidth with a low data rate, rectangular-like OOK signals discussed in the previous section have concentrated power spectrum that makes them vulnerable to frequency-selectivity of wireless fading. Alternative signal designs with more spread-out spectrum, but without compromising much on the ON/OFF energy split, can thus improve performance [^5] However, for the purpose of designing OOK synchronization signal, an OOK signal with narrow spectrum can provide benefits for performing frequency-synchronization in frequency-selective channels [30].. Below are methods to achieve this.